# Kai Zeng    Zhanqian Wu    Kaixin Xiong    Xiaobao Wei    Xiangyu Guo    Zhenxin Zhu    Kalok Ho    Lijun Zhou    Bohan Zeng    Ming Lu    Haiyang Sun    Bing Wang    Guang Chen    Hangjun Ye    Wentao Zhang [ [ [

### v3

@@ Abstract
Recent advancements in driving world models enable controllable generation of high-quality RGB videos or multimodal videos. Existing methods primarily focus on metrics related to generation quality and controllability. However, they often overlook the evaluation of downstream perception tasks, which are really crucial for the performance of autonomous driving. Existing methods usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). When we double the epochs in the baseline, the benefit of synthetic data becomes negligible. To thoroughly demonstrate the benefit of synthetic data, we introduce Dream4Drive, a novel synthetic data generation framework designed for enhancing the downstream perception tasks. Dream4Drive first decomposes the input video into several 3D-aware guidance maps and subsequently renders the 3D assets onto these guidance maps. Finally, the driving world model is fine-tuned to produce the edited, multi-view photorealistic videos, which can be used to train the downstream perception models. Dream4Drive enables unprecedented flexibility in generating multi-view corner cases at scale, significantly boosting corner case perception in autonomous driving. To facilitate future research, we also contribute a large-scale 3D asset dataset named DriveObj3D, covering the typical categories in driving scenarios and enabling diverse 3D-aware video editing. We conduct comprehensive experiments to show that Dream4Drive can effectively boost the performance of downstream perception models under various training epochs.
[…]
https://wm-research.github.io/Dream4Drive/
Figure 1: Dream4Drive demonstrates the effectiveness of synthetic data: with fewer than 2% synthetic samples, it consistently improves detection and tracking across epochs, outperforming previous data augmentation baselines under fair evaluation. 1× denotes the baseline training epoch; 2× and 3× mean doubling and tripling it.
@@ 1 Introduction
Perception tasks such as 3D object detection (li2022coda; li2024bevformer; wang2023exploring; wang2025survey) and 3D tracking (wang2023exploring; zhang20233d; han2025novel), which support planning and decision-making (jiang2023vad; hu2023planning), are extremely important in autonomous driving. The performance of perception models, however, is highly dependent on large-scale annotated training datasets (caesar2020nuscenes; wang2023we_asap). To ensure reliability in rare but critical safety scenarios, it is essential to gather adequate long-tail data. Although the autonomous driving community has developed a thorough 3D annotation pipeline to facilitate data acquisition (zhao2025survey), collecting long-tail data remains highly time-consuming and labor-intensive.
Driving world models based on diffusion model (LDM) and ControlNet (zhang2023adding) have been developed in autonomous driving to generate synthetic data from scene layouts (e.g., BEV maps, 3D bounding boxes) and text (gao2023magicdrive; wen2024panacea; guo2025genesis). However, they provide limited control over the poses and appearances of individual objects, which restricts the ability to generate diverse synthetic data (li2025realistic). Editing methods (singh2024genmm; liang2025driveeditor; yu2025vehiclesim) instead of generation extend driving world models by using reference images and 3D bounding boxes to modify object appearance and pose, allowing for diverse corner cases. However, single-view insertion limits their application in multi-view BEV perception. Reconstruction-based methods using NeRF or 3DGS enable precise control over geometric structures (chen2021geosim; zanjani2025gaussian). While geometry is preserved, sparse training views often lead to artifacts and incomplete renderings. Additionally, the lack of illumination modeling results in inconsistencies between inserted objects and the background.
More importantly, we argue that the data augmentation experiments of previous methods (wen2024panacea; li2024uniscene) are unfair, as they usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). We find that, under the same number of training epochs, large amounts of synthetic datasets offer little to no advantage and can even perform worse than using real data alone. As shown in Fig. 1, under the 2× epoch setting, models trained exclusively on real data achieve higher mAP and NDS compared to those trained on real and synthetic data. Given the importance of downstream perception tasks for autonomous driving, we believe it is essential to rethink the effectiveness of the driving world model as a synthetic data generator for these tasks.
@@
Specifically, Dream4Drive leverages a multi-view video inpainting model finetuned from the Diffusion Transformer (dit). Unlike prior methods that rely on sparse spatial controls (e.g., BEV maps and 3D bounding boxes), Dream4Drive uses dense 3D-aware guidance maps (yu2019inverserendernet; liang2025diffusion) (e.g., depth, normal, edge, cutout, and mask) to preserve the geometry and appearance of the original video, while editing them by rendering 3D assets into these maps. This design enables instance-level, cross-view consistent video editing, ensuring both visual realism and geometric fidelity. The generated videos not only achieve superior quality but can also be directly used to train state-of-the-art perception models.
To facilitate diverse 3D-aware video editing, we design a pipeline that automatically acquires high-quality 3D assets based on a target scene image or video of a desired asset category. We first apply the image segmentation model (ren2024grounded; lin2024draw) to localize and crop objects of the specified category, then employ the image generation model (wu2025qwenimagetechnicalreport) to generate multi-view consistent images of the target object. These images are fed into a mesh generation model (hunyuan3d2025hunyuan3d) to generate high-quality 3D assets. We present DriveObj3D, a large-scale 3D asset dataset that encompasses typical categories found in driving scenarios, to support future research. Our main contributions are:
@@ 2 Related Work
Video Generation in Autonomous Driving. High-quality data is crucial for training perception models in autonomous driving, motivating growing interest in driving video generation. Early approaches (yang2023bevcontrol; gao2023magicdrive; luo2025drvhierarchicalperceptiontemporalcognitionframework; li2024survey) employ diffusion models with ControlNet, conditioned on BEV maps (ma2024vision) and 3D bounding boxes, to generate paired image data (an2024mc; an2025concept; luo2024llm; luo2024panosent; luo2024nus). Recent works (gao2024magicdrivedit; jiang2024dive; li2024uniscene; ji2025cogen; guo2025genesis) adopt powerful Diffusion Transformers (DiTs) to further enhance generation quality. SubjectDrive (huang2024subjectdrive) uses an external subject bank to enhance vehicle appearance diversity. However, these methods depend on original scene layouts, which limit geometric diversity and struggle to generate high-quality long-tail corner cases, thereby restricting their effectiveness for downstream perception.
Video Editing in Autonomous Driving. To enrich scene diversity, object-level editing methods insert new objects into existing videos using reference images and 3D bounding boxes (singh2024genmm; liang2025driveeditor; buburuzan2025mobi; an2025unictokens). More recent NeRF- (mildenhall2021nerf) and 3DGS-based (kerbl20233dgs) approaches (chen2021geosim; huang2024s3g; chen2024omnire; wei2024emd) improve geometric fidelity, yet remain limited by sparse views, inconsistent lighting, and restricted background diversity. In contrast, our approach builds on generative models and introduces a novel 3D-aware video editing mechanism, enabling seamless insertion of diverse 3D assets and producing geometrically and visually diverse data that effectively boosts downstream performance.
@@ 3 Dream4Drive
Our goal is to generate high-quality synthetic videos from real videos with 3D box annotations and target 3D assets for training perception models. We first introduce the preliminaries in Sec. 3.1, then explain how to conduct 3D-aware scene editing in Sec. 3.2, and finally describe how to render the edited video from the guidance maps in Sec. 3.3. The overall framework is shown in Fig. 2.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs) (LDM) address the high computational cost of diffusion models by operating in a lower-dimensional latent space. Given an image $\mathbf{x}$, an encoder $\mathcal{E}$ is used to obtain the corresponding latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The forward diffusion process in the latent space is defined as a gradual noising process:
@@
ControlNets (controlnet2023) extend LDMs by incorporating additional conditioning signals $\mathbf{c}$ to provide finer control over the generation process. In particular, the reverse diffusion process is modified to condition on $\mathbf{c}$:
@@
For an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we use Depth Anything (yang2024depth) to obtain the depth map $D\in\mathbb{R}^{H\times W\times 1}$. The normal map $N\in\mathbb{R}^{H\times W\times 3}$ is then derived from the depth map. We also use OpenCV’s Canny edge detector to get the edge map $E\in\mathbb{R}^{H\times W\times 1}$. It is important to note that the depth, normal, and edge information within the foreground object regions are masked since our model needs to learn to generate an edited video based on the target 3D asset.
@@
Once we have obtained the depth map, normal map, and edge map for the background, as well as the rendered object image and object mask for the target 3D asset, we utilize a fine-tuned driving world model to render the edited video based on these 3D-aware guidance maps. This 3D-aware scene editing pipeline effectively utilizes the accurate pose, geometry, and texture information provided by 3D assets, ensuring geometric consistency in the results generated. Notably, our method does not depend on 3D bounding box embeddings for controlling object placement. Instead, we directly edit in 3D space, offering a more intuitive and reliable way to manage control.
@@
The video generation process in recent methods (wen2024panacea; li2024uniscene; guo2025genesis) relies on sparse spatial controls, such as bird’s eye view (BEV) maps and projected 3D bounding boxes. While this approach allows for some control over the synthesized content, it often falls short in achieving high-quality generation. Instead of relying on sparse spatial controls, we fine-tune a driving world model to generate edited videos from the dense 3D-aware guidance maps, including depth map ($D$), normal map ($N$), and edge map ($E$) for the background and the image ($O$) and mask ($M$) for the foreground object.
@@
We adopt rectified flow (liu2022flow) for stable sampling and classifier-free guidance (ho2022classifier) to balance text with multiple 3D geometric conditions, thereby improving controllability. The main training objective is a simplified diffusion loss that predicts the noise component at each timestep:
@@
where $\epsilon$ is Gaussian noise, $\epsilon_{\theta}$ the predicted noise, $\mathbf{z}_{t}$ the noisy latent at timestep $t$, and $\mathbf{c}$ all conditioning signals. To achieve precise instance-level control, we introduce a Foreground Mask Loss as ji2025cogen and an LPIPS loss (zhang2018unreasonable). The final training objective is:
@@ 4 DriveObj3D
To build a large-scale 3D assets for diverse 3D-aware video editing, we design a simple yet effective pipeline. The core idea is to decompose the asset generation process into three steps: (i) 2D instance segmentation; (ii) multi-view image generation; (iii) 3D mesh generation. As shown in Fig. 5, the pipeline inputs a video or image and the target asset category, and generates a 3D mesh for downstream applications.
Concretely, we first localize and segment the target object. Given an input image $I$ and category label class, Grounded-SAM (ren2024grounded) is applied to segment the object $I_{\text{target}}$. To overcome the occlusion of target object, we employ a multi-view image generation model Qwen-Image (wu2025qwenimagetechnicalreport). Conditioned on $I_{\text{target}}$, Qwen-Image synthesizes a set of novel views $\{I_{v}\}_{v=1}^{N}$, which are subsequently fed into a multi-view reconstruction model Hunyuan3D (hunyuan3d2025hunyuan3d) to recover the final 3D mesh.
Figure 6: Comparison of 3D asset generation across different methods. Our simple yet effective method produces better 3D assets across diverse categories in autonomous driving, outperforming existing baselines. Con_vehicle is construction vehicle; Pdes is Pedestrian; T_cone is traffic cone.
As shown in Fig. 6, assets generated by Text-to-3D methods (xiang2025structured) often exhibit style inconsistencies with the original data, while single-view approaches (hunyuan3d2025hunyuan3d) tend to produce incomplete assets. In contrast, our simple yet effective pipeline leverages multi-view synthesis to generate complete and high-fidelity assets, even under severe occlusions. To support large-scale downstream driving tasks, we construct a diverse asset dataset, DriveObj3D, covering a wide range of categories in driving scenarios in Fig. 7. All assets will be released publicly.
@@ 5.1 Setup
We utilize the nuScenes dataset (caesar2020nuscenes) for building 3D assets and finetuning our video generation model. It comprises 1000 scenes in total, with 700 designated for training, 150 for validation, and 150 for testing. Each scene contains a 20-second multi-view video captured by 6 cameras. More details on inserted scene and asset selection are given in the Appx. A.
@@
Following Panacea (wen2024panacea) and SubjectDrive (huang2024subjectdrive), we primarily evaluate how the generated data improves perceptual model performance on detection and tracking tasks. Detection metrics include nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Tracking metrics include Average MultiObject Tracking Accuracy (AMOTA), Precision (AMOTP), Recall, and Accuracy (MOTA).

### v4

# Rethinking Driving World Model as Synthetic Data Generator for Perception Tasks

@@ Abstract
Recent advancements in driving world models enable controllable generation of high-quality RGB videos or multimodal videos. Existing methods primarily focus on metrics related to generation quality and controllability. However, they often overlook the evaluation of downstream perception tasks, which are really crucial for the performance of autonomous driving. Existing methods usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). When we double the epochs in the baseline, the benefit of synthetic data becomes negligible. To thoroughly demonstrate the benefit of synthetic data, we introduce Dream4Drive, a novel synthetic data generation framework designed for enhancing the downstream perception tasks. Dream4Drive first decomposes the input video into several 3D-aware guidance maps and subsequently renders the 3D assets onto these guidance maps. Finally, the driving world model is fine-tuned to produce the edited, multi-view photorealistic videos, which can be used to train the downstream perception models. Dream4Drive enables unprecedented flexibility in generating multi-view corner cases at scale, significantly boosting corner case perception in autonomous driving. To facilitate future research, we also contribute a large-scale 3D asset dataset named DriveObj3D, covering the typical categories in driving scenarios and enabling diverse 3D-aware video editing. We conduct comprehensive experiments to show that Dream4Drive can effectively boost the performance of downstream perception models under various training epochs. Project website: https://wm-research.github.io/Dream4Drive/.
Figure 1: Dream4Drive demonstrates the effectiveness of synthetic data: with fewer than 2% synthetic samples, it consistently improves detection and tracking across epochs, outperforming previous data augmentation baselines under fair evaluation. 1× denotes the baseline training epochs; 2× and 3× represent twofold and threefold increases, respectively.
@@ 1 Introduction
Perception tasks such as 3D object detection (Li et al., 2022; 2024b; Wang et al., 2023a; 2025) and 3D tracking (Wang et al., 2023a; Zhang et al., 2023c; Han et al., 2025; Li et al., 2025b), which support planning and decision-making (Jiang et al., 2023; Hu et al., 2023; Shan et al., 2025; Hao et al., 2024), are extremely important in autonomous driving. The performance of perception models, however, is highly dependent on large-scale annotated training datasets (Caesar et al., 2020; Wang et al., 2023b). To ensure reliability in rare but critical safety scenarios, it is essential to gather adequate long-tail data. Although the autonomous driving community has developed a thorough 3D annotation pipeline to facilitate data acquisition (Zhao et al., 2025b), collecting long-tail data remains highly time-consuming and labor-intensive.
Driving world models based on diffusion and ControlNet (Rombach et al., 2022; Zhang et al., 2023b) generate synthetic data from scene layouts and text (Gao et al., 2023; Wen et al., 2024; Guo et al., 2025), but offer limited control over object pose and appearance, reducing data diversity (Li et al., 2025a). Editing-based methods (Singh et al., 2024; Liang et al., 2025b; Yu et al., 2025) improve this by inserting objects using reference images and 3D boxes, yet their single-view nature restricts use in multi-view BEV perception. Reconstruction-based approaches (NeRF, 3DGS) (Chen et al., 2021; Zanjani et al., 2025) provide geometric control but suffer from artifacts due to sparse views and lack illumination modeling, causing inconsistencies between inserted objects and the background.
More importantly, we argue that the data augmentation experiments of previous methods (Wen et al., 2024; Li et al., 2024a) are unfair, as they usually leverage a training strategy that first pretrains on synthetic data and finetunes on real data, resulting in twice the epochs compared to the baseline (real data only). We find that, under the same number of training epochs, large amounts of synthetic datasets offer little to no advantage and can even perform worse than using real data alone. As shown in Fig. 1, under the 2× epoch setting, models trained exclusively on real data achieve higher mAP and NDS compared to those trained on real and synthetic data. Given the importance of downstream perception tasks for autonomous driving, we believe it is essential to rethink the effectiveness of the driving world model as a synthetic data generator for these tasks.
@@
Specifically, Dream4Drive leverages a multi-view video inpainting model finetuned from the Diffusion Transformer (Peebles and Xie, 2023). Unlike prior methods that rely on sparse spatial controls (e.g., BEV maps and 3D bounding boxes), Dream4Drive uses dense 3D-aware guidance maps (Yu and Smith, 2019; Liang et al., 2025a) (e.g., depth, normal, edge, cutout, and mask) to preserve the geometry and appearance of the original video, while editing them by rendering 3D assets into these maps. This design enables instance-level, cross-view consistent video editing, ensuring both visual realism and geometric fidelity. The generated videos not only achieve superior quality but can also be directly used to train state-of-the-art perception models (Wang et al., 2023a).
To facilitate diverse 3D-aware video editing, we design a pipeline that automatically acquires high-quality 3D assets based on a target scene image or video of a desired asset category. We first apply the image segmentation model (Ren et al., 2024; Lin et al., 2024) to localize and crop objects of the specified category, then employ the image generation model (Wu et al., 2025) to generate multi-view consistent images of the target object. These images are fed into a mesh generation model (Hunyuan3D et al., 2025) to generate high-quality 3D assets. We present DriveObj3D, a large-scale 3D asset dataset that encompasses typical categories found in driving scenarios, to support future research. Our main contributions are:
@@ 2 Related Work
Video Generation in Autonomous Driving. High-quality data is crucial for training perception models in autonomous driving, motivating growing interest in driving video generation. Early approaches (Yang et al., 2023; Gao et al., 2023; Wen et al., 2024) employ diffusion models with ControlNet, conditioned on BEV maps (Ma et al., 2024b) and 3D bounding boxes, to generate paired image data (Zhou et al., 2024; An et al., 2024; 2025b). Recent works (Gao et al., 2024a; Jiang et al., 2024; Li et al., 2024a; Ji et al., 2025; An et al., 2026; Li et al., 2025c) adopt powerful Diffusion Transformers (DiTs) to further enhance generation quality. SubjectDrive (Huang et al., 2024a) uses an external subject bank to enhance vehicle appearance diversity. However, these methods depend on original scene layouts, which limit geometric diversity and struggle to generate high-quality long-tail corner cases, thereby restricting their effectiveness for downstream perception.
Video Editing in Autonomous Driving. To enrich scene diversity, object-level editing methods insert new objects into existing videos using reference images and 3D bounding boxes (Singh et al., 2024; Liang et al., 2025b; Buburuzan et al., 2025; An et al., 2025a). More recent NeRF- (Mildenhall et al., 2021) and 3DGS-based (Kerbl et al., 2023) approaches (Chen et al., 2021; Huang et al., 2024b; Chen et al., 2024; Wei et al., 2024; Zhu et al., 2025) improve geometric fidelity, yet remain limited by sparse views, inconsistent lighting, and restricted background diversity. In contrast, our approach builds on generative models and introduces a novel 3D-aware video editing mechanism, enabling seamless insertion of diverse 3D assets and producing geometrically and visually diverse data that effectively boosts downstream performance.
@@ 3 Dream4Drive
Our goal is to generate high-quality synthetic videos based on real video with ground-truth 3D box annotations and a target 3D asset, creating synthetic data for training downstream perception models. We first introduce the preliminaries in Sec. 3.1, then explain how to conduct 3D-aware scene editing in Sec. 3.2, and finally describe how to render the edited video from the guidance maps in Sec. 3.3. The overall framework is shown in Fig. 2.
@@ 3.1 Preliminaries
Latent Diffusion Models (LDMs) (Rombach et al., 2022) address the high computational cost of diffusion models by operating in a lower-dimensional latent space. Given an image $\mathbf{x}$, an encoder $\mathcal{E}$ is used to obtain the corresponding latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The forward diffusion process in the latent space is defined as a gradual noising process:
@@
ControlNets (Zhang et al., 2023a) extend LDMs by incorporating additional conditioning signals $\mathbf{c}$ to provide finer control over the generation process. In particular, the reverse diffusion process is modified to condition on $\mathbf{c}$:
@@
To overcome the limitations of implicit 3D controls (like bounding boxes or BEV maps) in precisely placing assets, we introduce a new form of 3D-aware guidance maps.
For an input RGB image $I\in\mathbb{R}^{H\times W\times 3}$, we use Depth Anything (Yang et al., 2024) to obtain the depth map $D\in\mathbb{R}^{H\times W\times 1}$. The normal map $N\in\mathbb{R}^{H\times W\times 3}$ is then derived from the depth map. We also use OpenCV’s Canny edge detector to get the edge map $E\in\mathbb{R}^{H\times W\times 1}$. It is important to note that the depth, normal, and edge information within the foreground object regions are masked since our model needs to learn to generate an edited video based on the target 3D asset.
@@
Once we have obtained the depth map, normal map, and edge map for the background, as well as the rendered object image and object mask for the target 3D asset, we utilize a fine-tuned driving world model to render the edited video based on the 3D-aware guidance maps. This 3D-aware scene editing pipeline effectively utilizes the accurate pose, geometry, and texture information provided by 3D assets, ensuring geometric consistency in the results generated. Notably, our method does not depend on 3D bounding box embeddings for controlling object placement. Instead, we directly edit in 3D space, offering a more intuitive and reliable way to manage control.
@@
The video generation process in recent methods (Wen et al., 2024; Li et al., 2024a; Guo et al., 2025) relies on sparse spatial controls, such as bird’s eye view (BEV) maps and projected 3D bounding boxes. While this approach allows for some control over the synthesized content, it often falls short in achieving high-quality generation. Instead of relying on sparse spatial controls, we fine-tune a driving world model to generate edited videos from the dense 3D-aware guidance maps, including depth map ($D$), normal map ($N$), and edge map ($E$) for the background and the image ($O$) and mask ($M$) for the foreground object.
@@
We adopt rectified flow (Liu et al., 2022) for stable sampling and classifier-free guidance (Ho and Salimans, 2022) to balance text with multiple 3D geometric conditions, thereby improving controllability. The main training objective is a simplified diffusion loss that predicts the noise component at each timestep:
@@
where $\epsilon$ is Gaussian noise, $\epsilon_{\theta}$ the predicted noise, $\mathbf{z}_{t}$ the noisy latent at timestep $t$, and $\mathbf{c}$ all conditioning signals. To achieve precise instance-level control, we introduce a Foreground Mask Loss as Ji et al. (2025) and an LPIPS loss (Zhang et al., 2018). The final training objective is:
@@ 4 DriveObj3D
The diversity of the synthesized scenes is directly determined by the richness of the asset library used for insertion. To build a large-scale 3D assets for diverse 3D-aware video editing, we design a simple yet effective pipeline. The core idea is to decompose the asset generation process into three steps: (i) 2D instance segmentation; (ii) multi-view image generation; (iii) 3D mesh generation. As shown in Fig. 5, the pipeline inputs a video or image and the target asset category, and generates a 3D mesh for downstream applications.
Concretely, we first localize and segment the target object. Given an input image $I$ and category label class, Grounded-SAM (Ren et al., 2024) is applied to segment the object $I_{\text{target}}$. To overcome the occlusion of target object, we employ a multi-view image generation model Qwen-Image (Wu et al., 2025). Conditioned on $I_{\text{target}}$, Qwen-Image synthesizes a set of novel views $\{I_{v}\}_{v=1}^{N}$, which are subsequently fed into a multi-view reconstruction model Hunyuan3D (Hunyuan3D et al., 2025) to recover the final 3D mesh.
Figure 6: Comparison of 3D asset generation across different methods. Our simple yet effective method produces better 3D assets across diverse categories in autonomous driving, outperforming existing baselines. Con_vehicle is construction vehicle; Pdes is Pedestrian; T_cone is traffic_cone.
As shown in Fig. 6, assets generated by Text-to-3D methods (Xiang et al., 2025) often exhibit style inconsistencies with the original data, while single-view approaches (Hunyuan3D et al., 2025) tend to produce incomplete assets. In contrast, our simple yet effective pipeline leverages multi-view synthesis to generate complete and high-fidelity assets, even under severe occlusions. To support large-scale downstream driving tasks, we construct a diverse asset dataset, DriveObj3D, covering a wide range of categories in driving scenarios in Fig. 7. All assets will be released publicly.
@@ 5.1 Setup
We utilize the nuScenes dataset (Caesar et al., 2020) for building 3D assets and finetuning our video generation model. It comprises 1000 scenes in total, with 700 designated for training, 150 for validation, and 150 for testing. Each scene contains a 20-second multi-view video captured by 6 cameras. More details on inserted scene and asset selection are given in the Appx. A.
1x Epochs 2x Epochs
(28130) (+420, <2%) (28130) (Wang et al., 2024a) (Lu et al., 2024) (Gao et al., 2023) (Wen et al., 2024) (Huang et al., 2024a) (Ours)
mAP $\uparrow$ 34.5 36.1 38.4 35.8 36.2 35.4 37.1 38.1 38.7
mAVE $\downarrow$ 29.1 28.9 27.7 123.4 27.3 26.4 26.8
NDS $\uparrow$ 46.9 47.8 50.4 39.5 18.1 39.8 49.2 50.2 50.6
Table 1: Comparison of detection under different training epochs. * indicates the evaluation of WoVoGen is only on the vehicle classes of cars, trucks, and buses.
1x Epochs 2x Epochs
Real Dream4Drive Real Panacea SubjectDrive Dream4Drive
(28130) (+420, <2%) (28130) (Wen et al., 2024) (Huang et al., 2024a) (Ours)
AMOTA $\uparrow$ 30.1 31.2 34.1 33.7 33.7 34.4
AMOTP $\downarrow$ 137.9 135.4 134.1 135.3 135.3 133.5
Table 2: Comparison of tracking under different training epochs.
@@
Following Panacea (Wen et al., 2024) and SubjectDrive (Huang et al., 2024a), we primarily evaluate how the generated data improves perceptual model performance on detection and tracking tasks. Detection metrics include nuScenes Detection Score (NDS), mean Average Precision (mAP), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Tracking metrics include Average MultiObject Tracking Accuracy (AMOTA), Precision (AMOTP), Recall, and Accuracy (MOTA).